### v1

@@ 5 Numerical experiments
We first verify that the linear systems encountered in the numerical solution of the inpainting bilevel problem are indeed similar. We consider the sequence of the Hessians , right hand sides and solution that were determined without recycling, and compute the relative difference of the Frobenius norm for each of these 3 sequences. In Figure 3 we plot the relative differences against the index of the linear system and we see that the sequence of linear systems are similar with respect to all of these quantities, with a relative differences of order being observed in all instances. Thus, Krylov subspace recycling may be a viable approach to reduce the overall computational cost.
@@
Projected vs full dimensional matrix decompositions
We have motivated Ritz vectors and Ritz generalized singular vectors as approximations of eigenvectors and generalized singular vectors respectively, restricted to some subspace Computing the full-dimensional matrix decompositions (i.e. taking in either (20) or (24)) is not viable in practice due to the high computational cost. However, to provide insight into how the projected approximations compare to them, we solve the sequence of systems with Ritz-S, RGen-L(R), and the corresponding full-dimensional matrix decomposition strategies, Eig-S and GSVD-L(R), where we still select the recycle space to have dimension at most .
In Figure 5 we plot the obtained number of iterations to terminate RMINRES, cumulative number of RMINRES iterations, and the relative error of the obtained hypergradients. We see that Ritz-S and RGen-L(R) regularly terminate earlier than their full-dimensional matrix decomposition counterparts, Eig-S and GSVD-L(R), respectively. Since the residual norm stopping criterion is used in all of these strategies, this earlier termination of Ritz-S and RGen-L(R) suggests that performing a projection can compact useful information of the solution space from the full-dimensional matrix decomposition. We see that GSVD-L(R) yields hypergradients an order of magnitude more accurate than the other methods, which obtain similar quality hypergradients to one-another. This supports the claim that encoding information of both the Hessian and adjoint Jacobian in the recycle space is a valid strategy to yield high quality hypergradients.
@@
We plot the performance of the considered recycling strategies in Figure 6. We see that Ritz vectors associated with the smallest Ritz values (i.e., Ritz-S) yield the largest reduction in iterations, with those associated with large Ritz values achieving little improvement over not employing recycling at all. For harmonic Ritz vectors, information associated with the largest harmonic Ritz values results in the largest decrease of iterations, comparable to Ritz-S. Similarly, we see that Ritz generalized singular vectors associated with the largest Ritz generalized singular values achieve the largest reduction in iterations. Given the performance of Ritz vectors this is to be expected. Indeed, should the Ritz generalized singular values are the reciprocals of the Ritz values. Furthermore, using right rather than left Ritz generalized singular vectors yields a larger reduction in number of iterations. Utilizing the new stopping criterion (28) roughly halves the total number of iterations compared to using the residual norm. Indeed, in the bottom row of Figure 6 we see that although Ritz vectors perform better than Ritz generalized vectors when using the residual norm, in employing the new stopping criterion, right Ritz generalized vectors yield a greater reduction in number of iterations. We see that the new stopping criterion often results in a larger relative error in the associated hypergradient compared to other recycling strategies and stopping criteria. It should be noted that the methods which do not employ the new stopping criterion do not have a way to evaluate (an approximation of) the hypergradient error and so the relative error is unknown in practice. In contrast, the strategies that employ the new stopping criterion terminate when an approximation of the hypergradient error is which in this example roughly coincides with the magnitude of the hypergradient relative error.
@@
We compare the cumulative number of iterations in Figure 6, where the residual norm stopping criterion was used, to Figure 7, where results for using the true hypergradient error as the stopping criterion are displayed. We see that no recycling takes 1447 total iterations, which is little reduction compared to using a residual norm as a stopping criterion which required 1500 iterations. This indicates that, in building the Krylov subspace from scratch, all iterations of MINRES are required to generate a solution space that can lead to a good hypergradient. In contrast, by recycling information from the previous solution space, Ritz-S reduces from 764 total iterations to 652 (81.9%), and RGen-L(R) reduces from 871 to 713 (85.3%). This is to be expected since an accurate solution to the linear system is only a proxy for an accurate hypergradient. Interestingly, in Figure 6 we see that RGen-L(R)-NSC, which employed the approximation (28) of the hypergradient error as the stopping criterion, took 500 iterations. This indicates that the proposed approximation (28) often acts as a lower bound of the true hypergradient error (27). We explore this aspect more thoroughly in the next experiment.
@@
To understand how good of an approximation our proposed projected hypergradient error (28) is of the true hypergradient error (27), we solve the sequence of systems with RGen-L(R)-NSC, i.e. we employ the approximated hypergradient error (28) as the stopping criterion, and study the iterates from each RMINRES call with respect to quantities relevant to the proposed approximation, of which we give details below.
Recall that the approximated hypergradient error (28) requires computing the GSVD involving projected onto a subspace spanned by that is, a subspace consisting of information related to Thus, errors in the approximation can occur not only in the fact that a projection is being performed, but also by potential suboptimality of the subspace we are projecting onto. Since we solve the sequence of Hessians with RGen-L(R)-NSC, we can evaluate the approximated hypergradient error that utilizes the GSVD of projected onto that is, information associated with the current linear system, rather than the previous one. Although this is only possible if the entire sequence is solved in advance, in this way, we can isolate the effect of both performing a projection, and performing a projection onto a subspace related to the previous system.
In Figure 8 we plot evaluations of the true (unprojected) hypergradient error, our proposed projected approximation that uses information of the previous system, and the approximation that utilizes information of the current system. The proposed approximation of the hypergradient error often serves as a lower bound for the other quantities, with the true hypergradient error often being an upper bound for the projected approximations.
Next, we select a few different Hessian systems and consider how the different quantities change throughout each RMINRES run, and plot the results in Figure 9. We see that projection onto a subspace with information associated with the current linear system serves as a good approximation of the true hypergradient error, especially for early iterations. This could motivate using a strategy inspired by [62], in which a recycle space to be employed for the next linear system solve is updated during the RMINRES run with information of the current system’s solution space. In particular, one could consider a stopping criterion that employs this iteratively updated recycle space. Similar to Figure 8, the proposed hypergradient error approximation regularly serves as a lower bound for the other quantities.
@@
To construct the th recycle space which will be used in solving the linear system involving we currently consider a matrix decomposition (be it the eigendecomposition or GSVD) which involves where requires that we store the entire Krylov subspace basis between linear system solves. In [62] it is proposed to consider a matrix decomposition concerning as then one can avoid the need to store the whole basis and exploit matrix decompositions involving and ; see [62] for details. We investigate the influence these two choices have on the recycling strategy. We refer to a recycling strategy that employs the matrix proposed in [62], that is, utilizing as an inner recycling strategy, since the recycle space for the next linear system is determined inside the RMINRES run for the current linear system. Similarly, we refer to the choice that we primarily employ, utilizing as an outer recycling strategy.
@@
We now investigate how the dimension of the recycling space affects performance. Indeed, to both construct Krylov vectors that are orthogonal to the recycle space, as in the augmented Lanczos relation (16), and update the part of the solution that lies within the recycle space, extra FLOPS are required for a single iteration of RMINRES versus standard MINRES; see Appendix B. Additionally, the cost of the matrix decomposition employed to determine the Ritz-type vectors will depend on the dimension of the recycle space. Thus, one must balance the reduction of iterations to the computational overhead associated with a given recycling space dimension.
@@ 6 Conclusions and outlook
Future work includes a rigorous analysis of both optimality of Ritz generalized singular vectors and the accuracy of the hypergradient error approximation. While extensive numerical experiments verify the proposed recycling strategy as a proof of concept, an implementation in a spirit similar to [62] is to be determined, wherein the Lanczos relation is exploited within the GSVD computation. Large-scale applications remain to be explored, specifically, those in which solutions of linear systems are employed in a matrix-vector product, for example, differentiable architecture search [39]. Using the proposed recycling strategy in conjunction with other computational saving techniques, such as inexact methods [15, 16, 49, 53], should be investigated, with special attention given to potential interaction regarding the hypergradient error approximation.
@@ Appendix A Calculating the residual vector for RMINRES
While MINRES provides an efficient way to determine the residual norm, to employ the alternative stopping criterion discussed in Section 4 we require access to the residual vector explicitly. Towards this, notice that it from Algorithm 3 line 24 the residual vector will satisfy
@@
and we already compute in line 10 for the construction of Since is already stored for RMINRES, we need only store two additional vectors, and to determine
@@ Appendix B Computational cost of RMINRES
We remark that, due to the additional 3 term recurrence that is employed to update the solution with respect to the recycle space, each iteration of RMINRES has an additional cost independent of the recycle dimension. Finally, (31) is only the cost associated with a run of RMINRES and neglects the cost to determine the recycle space itself.

### v2

@@ 5.1 MNIST Inpainting
We first experimentally verify that the linear systems encountered in the numerical solution of the inpainting bilevel problem are indeed similar. We consider the sequence of the Hessians , right hand sides and solutions that were determined without recycling, and compute the relative difference in the Frobenius norm for each of these 3 sequences. In Figure 3 we plot such quantities and we eventually see relative differences of order in all instances. Thus, Krylov subspace recycling may be a viable approach to reduce the overall computational cost. We emphasise that, in general, although outside the scope of this paper, such experimental validation may be supported by theoretical derivations, exploiting known properties of the Hessians (linked to the properties of the matrix and the filters ), and how these are affected by multiplications by , leading to an analysis similar to the one in [40]. Specifically, in [40], for a sequence of linear systems associated to the discretization of differential operators, and for consecutive systems that differ by a low enough rank plus small enough norm perturbation, the authors conclude that approximate invariant subspaces (good for recycling) are given by the (approximate) eigenvectors corresponding to the (approximate) small eigenvalues, that appear in disjoint enough clusters.
@@
Approximated versus full dimensional matrix decompositions
We have motivated Ritz vectors and Ritz generalized singular vectors as approximations of eigenvectors and generalized singular vectors respectively, restricted to some subspace Computing the full-dimensional matrix decompositions (i.e. taking in either (20) or (24)) is not viable in practice due to the high computational cost. However, to provide insight into how the approximations compare to them, we solve the sequence of systems with Ritz-S, RGen-L(R), and the corresponding full-dimensional matrix decomposition strategies, Eig-S and GSVD-L(R), where we still select the recycle space to have dimension at most .
In Figure 5 we plot the number of iterations to terminate RMINRES for each system, the cumulative number of RMINRES iterations, the relative error of the obtained hypergradients, and the associated upper level functional. We see that Ritz-S and RGen-L(R) regularly terminate earlier than their full-dimensional matrix decomposition counterparts, Eig-S and GSVD-L(R), respectively. Since the residual norm stopping criterion is used in all of these strategies, this earlier termination of Ritz-S and RGen-L(R) suggests that considering can compact useful information of the solution space from the full-dimensional matrix decomposition. Indeed, Ritz vectors can be viewed as projected approximations of eigenvectors. We see that GSVD-L(R) yields hypergradients an order of magnitude more accurate than the other methods, which obtain similar quality hypergradients to one-another. This supports the claim that encoding information of both the Hessian and adjoint Jacobian in the recycle space is a valid strategy to yield high quality hypergradients.
@@
We plot the performance of the considered recycling strategies in Figure 6. We see that Ritz vectors associated with the smallest Ritz values (i.e., Ritz-S) yield the largest reduction in iterations, with those associated with large Ritz values achieving little improvement over not employing recycling at all. For harmonic Ritz vectors, information associated with the largest harmonic Ritz values results in the largest decrease of iterations, comparable to Ritz-S. Similarly, we see that Ritz generalized singular vectors associated with the largest Ritz generalized singular values achieve the largest reduction in iterations. Given the performance of Ritz vectors this is to be expected. Indeed, should the Ritz generalized singular values are the reciprocals of the Ritz values. Furthermore, using right rather than left Ritz generalized singular vectors yields a larger reduction in number of iterations. Utilizing the new stopping criterion (28) roughly halves the total number of iterations compared to using the residual norm. Indeed, in the bottom row of Figure 6 we see that although Ritz vectors perform better than Ritz generalized vectors when using the residual norm, in employing the new stopping criterion, right Ritz generalized vectors yield a greater reduction in number of iterations. We see that the new stopping criterion often results in a larger relative error in the associated hypergradient compared to other recycling strategies and stopping criteria. It should be noted that the methods which do not employ the new stopping criterion do not have a way to evaluate (an approximation of) the hypergradient error, which is unknown in practice. In contrast, the strategies that employ the new stopping criterion terminate when an approximation of the hypergradient error is which, in this example, roughly coincides with the magnitude of the hypergradient relative error.
@@
We display the results using the true hypergradient error as the stopping criterion inFigure 7. We see that no recycling takes 1447 total iterations, which is little reduction compared to using a residual norm as a stopping criterion which required 1500 iterations (see Figure 6). This indicates that, in building the Krylov subspace from scratch, all iterations of MINRES are required to generate a solution space that can lead to a good hypergradient. In contrast, by recycling information from the previous solution space, Ritz-S and RGen-L(R) reduce from 764 and 871 total when using the residual norm (see Figure 6) to 652 (81.9%) and 713 (85.3%), respectively, when using the true hyptergradient error. This is to be expected since an accurate solution to the linear system is only a proxy for an accurate hypergradient. Interestingly, in Figure 6 we see that RGen-L(R)-NSC, which employed the approximation (28) of the hypergradient error as the stopping criterion, took 500 iterations. This indicates that the proposed approximation (28) often acts as a lower bound of the true hypergradient error (27). We explore this aspect more thoroughly in the next experiment.
@@
To understand how good our proposed approximation of the hypergradient error (28) is, with respect to the true hypergradient error (27), we solve the sequence of systems with RGen-L(R)-NSC, i.e. we employ the approximated hypergradient error (28) as the stopping criterion, and study the iterates from each RMINRES call with respect to quantities relevant to the proposed approximation, of which we give details below.
Recall that the approximated hypergradient error (28) requires computing the GSVD involving a projection of onto a subspace consisting of information related to namely, Thus, errors in the approximation can occur not only in the fact that , but also by potential suboptimality of the subspace spanned by
Since we solve the sequence of Hessians with RGen-L(R)-NSC, we can evaluate the approximated hypergradient error that utilizes the GSVD of projected onto that is, information associated with the current linear system, rather than the previous one. Although this is only possible if the entire sequence is solved in advance, in this way, we can isolate the effect of both performing a projection, and performing a projection onto a subspace related to the previous system.
In Figure 8 we plot evaluations of the true hypergradient error, our proposed approximation that uses information of the previous system, and the approximation that utilizes information of the current system. The proposed approximation of the hypergradient error often serves as a lower bound for the other quantities, with the true hypergradient error often being an upper bound for the approximations.
Next, we select a few different Hessian systems and consider how the different quantities change throughout each RMINRES run, and plot the results in Figure 9. We see that projection onto a subspace with information associated with the current linear system serves as a good approximation of the true hypergradient error, especially for early iterations. This could motivate using a strategy inspired by [75], in which a recycle space to be employed for the next linear system solve is updated during the RMINRES run with information of the current system’s solution space. In particular, one could consider a stopping criterion that employs this iteratively updated recycle space. Similar to Figure 8, the proposed hypergradient error approximation regularly serves as a lower bound for the other quantities. Interestingly, for the 94th Hessian system one approximation is slightly larger than the true hypergradient error, which suggests that the proposed approximation may not be able to be formally viewed as a projected version.
@@
To construct the th recycle space which will be used in solving the linear system involving we currently consider a matrix decomposition (be it the eigendecomposition or GSVD) which involves where requires that we store the entire Krylov subspace basis between linear system solves. In [75] it is proposed to consider a matrix decomposition concerning as then one can avoid the need to store the whole basis and exploit matrix decompositions involving and Specifically, for each linear system, our implementation requires storage of an matrix, whereas [75] requires storage of a matrix, where typically is chosen in advance such that for all see [75] for details. We investigate the influence these two choices have on the recycling strategy. We refer to a recycling strategy that employs the matrix proposed in [75], that is, utilizing as an inner recycling strategy, since the recycle space for the next linear system is determined inside the RMINRES run for the current linear system. Similarly, we refer to the choice that we primarily employ, utilizing as an outer recycling strategy.
@@
We now investigate how the dimension of the recycling space affects performance. Indeed, to both construct Krylov vectors that are orthogonal to the recycle space, as in the augmented Lanczos relation (16), and update the part of the solution that lies within the recycle space, extra FLOPS are required for a single iteration of RMINRES versus standard MINRES; see supplementary material. Additionally, the cost of the matrix decomposition employed to determine the Ritz-type vectors will depend on the dimension of the recycle space. Thus, one must balance the reduction of iterations to the computational overhead associated with a given recycling space dimension.
@@
5.2 BSDS300 Deconvolution
We now consider an application that resembles a bilevel problem (30) that is solved in practice. For images of the BSDS300 dataset [49] we take random crops (so and ), which have been convolved with a Gaussian kernel of standard deviation and then corrupted by Gaussian noise of level We learn filters of size for the FoE regularizer (30) () with which is motivated in the original work of [64]. Due to the non-convexity of the Hessian of the lower level is no longer guaranteed to be positive definite - further justifying the choices of MINRES over CG for this application. While the expression for hypergradients (7) may also no longer be valid, due to potential non-invertability of the Hessian, we follow common practice and construct hypergradients following the construction of (9) regardless. We solve (30) without recycling using Adam for 50 epochs with a batch size of 16 and fixed steplength of For the lower-level we employ L-BFGS with backtracking (memory size 10). A stopping tolerance of is used for all involved optimizers and we allow for a maximum of 16000 iterations for both the lower level and Hessian system solvers. In line with [10, 41], the filters are initialized with the non-constant 2D discrete cosine transform basis filters. Although the training data is split across mini-batches, each individual sample will, due to the fixed steplength, only require a single lower level and Hessian system solve each epoch. We plot the accumulated timings of the lower level and Hessian system solvers across all training samples in Figure 12, where we see that solving the Hessian systems takes over 29% of computation time.
Reconstructions of on an image from the test when using the learned and initial filters, that is, the DCT basis, are shown in top row of Figure 13. The learned filters succesfuly recover a meaningful reconstruction. The initial and learned filters are displayed in the bottom row of Figure 13, where we observe that the low frequency filters have changed the most.
We re-solve the sequence of Hessians associated with every sample of training data data using the recycling strategies that performed best in Section 5.1, namely, Rits-S, RGen-L(R), and RGen-L(R)-NSC. The performance of said strategies for solving the sequence of systems associated with first sample of training data is displayed in the top row of Figure 14. Both Ritz-S and RGen-L(R) reduce the overall number of RMINRES iterations by approximately 27%, with Ritz-S slightly out-performing RGen-L(R). Employing the new stopping criterion (28) leads to a reduction of 39% while maintaining hypergradients of similar quality. The bottom row of Figure 14 shows the accumulated performance across all sequences of systems solved during training. Both Ritz-S and RGen-L(R) yield a reduction of while the RGen-L(R) obtains a larger reduction of All methods attain a smaller reduction than what was observed in Section 5.1. This is likely a result of using mini-batches in training, as potentially multiple parameter updates have been performed before a specific training sample is encountered again.
@@ 6 Conclusions and outlook
Future work includes a rigorous analysis of both optimality of Ritz generalized singular vectors and the accuracy of the hypergradient error approximation, as well as the study of alternative optimal approaches that take into account the premultiplication of in the hypergradient computations. While extensive numerical experiments verify the proposed recycling strategy as a proof of concept, an implementation in a spirit similar to [75] is to be determined, wherein the Lanczos relation is exploited within the GSVD computation. Large-scale applications remain to be explored, specifically, those in which solutions of linear systems are employed in a matrix-vector product, for example, differentiable architecture search [47]. Using the proposed recycling strategy in conjunction with other computational saving techniques, such as inexact methods [21, 22, 60, 66], should be investigated, with special attention given to potential interaction regarding the hypergradient error approximation.
@@ Appendix A Calculating the residual vector for RMINRES
While MINRES provides an efficient way to determine the residual norm, to employ the alternative stopping criterion discussed in Section 4 we require explicit access to the residual vector Towards this, notice that from Algorithm 4 line 24 the residual vector will satisfy
@@
and we already compute in line 10 for the construction of Since is already stored in RMINRES, we need only store two additional vectors, and to determine
@@ Appendix B Computational cost of RMINRES
We remark that, due to the additional 3 term recurrence that is employed to update the solution with respect to the recycle space, each iteration of RMINRES has an additional cost independent of the recycle dimension. Finally, (32) is only the cost associated with a run of RMINRES and neglects the cost to determine the recycle space itself. We remark that, depending on the application (that is, the value of and number of iterations RMINRES is expected to take, it is possible implement a more computationally effective algorithm than 4 [50].